OpenStudy (anonymous):

How to show 713.65 in expanded notation?

OpenStudy (anonymous):

It could be 71.365 x 10^-1 @steephpalaciosss

OpenStudy (anonymous):

Thanks for the help!😊

OpenStudy (anonymous):

No problem want to be a Fan and Give me a Medal???

OpenStudy (aum):

713.65 in expanded notation: \(713.65 = 7 \times 100 + 1 \times 10 + 3 \times 1 + 6 \times \frac{1}{10} + 5 \times \frac{1}{100}\)